# Symmetrical Flow Matching: Unified Image Generation, Segmentation, and Classification with Score-Based Generative Models

### v1

@@ Abstract
Flow Matching has emerged as a powerful framework for learning continuous transformations between distributions, enabling high-fidelity generative modeling. This work introduces Symmetrical Flow Matching (SymmFlow), a new formulation that unifies semantic segmentation, classification, and image generation within a single model. Using a symmetric learning objective, SymmFlow models forward and reverse transformations jointly, ensuring bi-directional consistency, while preserving sufficient entropy for generative diversity. A new training objective is introduced to explicitly retain semantic information across flows, featuring efficient sampling while preserving semantic structure, allowing for one-step segmentation and classification without iterative refinement. Unlike previous approaches that impose strict one-to-one mapping between masks and images, SymmFlow generalizes to flexible conditioning, supporting both pixel-level and image-level class labels. Experimental results on various benchmarks demonstrate that SymmFlow achieves state-of-the-art performance on semantic image synthesis, obtaining FID scores of 11.9 on CelebAMask-HQ and 7.0 on COCO-Stuff with only 25 inference steps. Additionally, it delivers competitive results on semantic segmentation and shows promising capabilities in classification tasks. The code will be publicly available.
Introduction
@@
Comprehending semantic content is a key challenge in computer vision. Classification (Krizhevsky, Sutskever, and Hinton 2012; He et al. 2016; Dosovitskiy et al. 2020; Liu et al. 2022) and segmentation (Long, Shelhamer, and Darrell 2015; Ronneberger, Fischer, and Brox 2015; Xie et al. 2021; Cheng et al. 2022) allow models to analyze and structure images, while generative modeling enables the synthesis of new content (Radford, Metz, and Chintala 2015; Zhu et al. 2017; Ho, Jain, and Abbeel 2020; Rombach et al. 2022). Ideally, a unified framework would bridge these tasks, allowing models to both interpret and generate images in a two-way manner. It is conceivable that the ability to accurately comprehend and disentangle visual structures facilitates the generation of more semantically coherent and visually realistic images. Conversely, strong generative capabilities may aid in learning more expressive representations of images, as generating plausible content requires an implicit understanding of object relationships, textures, and context (He et al. 2022). These forward and backward relationships suggest that advances in one direction could naturally benefit the other, which fuels the pursuit of models that integrate both understanding and synthesis within a single cohesive framework.
@@
Most existing vision-only approaches treat these tasks individually. For classification, the extracted features are used in fully connected layers to obtain probability predictions for each class, while for segmentation, the features are used to train decoders for dense predictions. On the other hand, generative models, such as generative adversarial networks (GANs) (Karras et al. 2021), diffusion models (Karras et al. 2022), score matching models (Song et al. 2020) and Flow Matching models (Lipman et al. 2022), synthesize images from a prior distribution. Although recent work has explored diffusion models for classification (Li et al. 2023a) and segmentation (Wu et al. 2024a, b), these adaptations introduce significant limitations: classification is slow due to the need for iterative sampling across all possible classes, and segmentation frameworks are restricted to generating masks, lacking the ability to map back to realistic images.
@@
Recent works, such as SemFlow (Wang et al. 2024b) and DepthFM (Gui et al. 2024), have sought to unify image generation with semantic segmentation and depth estimation within a single generative framework. However, these models still suffer from the following key limitations: (1) they do not perform classification, restricting their applicability; (2) the image quality remains inferior to that of purely generative models; and (3) the models enforce a strict one-to-one mapping between segmentation or depth masks and images, requiring them to have the same number of channels, which limits their flexibility.
@@
Related Work
Flow Matching
@@
Flow Matching (FM) (Lipman et al. 2022; Liu, Gong, and Liu 2022; Albergo and Vanden-Eijnden 2022) is a generative modeling framework that learns a velocity field to transform a source distribution into a target distribution through a continuous flow. By parameterizing this field with a neural network, the model becomes a Neural ODE (Chen et al. 2018), allowing efficient sampling via numerical integration. It has advanced state-of-the-art performance in diverse applications, including image (Esser et al. 2024), and audio generation (Vyas et al. 2023; Le et al. 2023), as well as protein modeling (Huguet et al. 2024) and robotics (Black et al. 2024). FM generalizes Continuous Normalizing Flows (CNFs) (Chen et al. 2018; Grathwohl et al. 2018) by removing the need for simulation during training, thereby making it computationally efficient. It also provides a unifying perspective on generative modeling, encompassing diffusion models, which can be seen as a special case where the probability path is defined via stochastic differential equations (Lipman et al. 2024).
Generative Classifiers
@@
Several seminal works (Hinton 2007; Ranzato et al. 2011) have emphasized the importance of modeling the data distribution to enhance discriminative feature learning. Early approaches trained deep belief networks (Hinton, Osindero, and Teh 2006) to encode image data as latent representations, which were then used for recognition tasks. More recent advances in generative modeling have demonstrated the ability to learn efficient representations for global prediction tasks (He et al. 2022; Croce, Castellucci, and Basili 2020). In addition, generative models have been shown to improve adversarial robustness and calibration (Huang et al. 2020). However, most prior work either jointly trains generative and discriminative models, or fine-tunes generative representations for downstream tasks. Diffusion Classifier (Li et al. 2023a) specifically investigates the effectiveness of using diffusion models as image classifiers, albeit with severe inference constraints.
Semantic Segmentation
@@
Semantic segmentation aims to assign a semantic label to each pixel in an image. Conventional approaches rely on discriminative models, combining a strong feature extraction backbone with a task-specific decoder head for mask prediction (Cheng, Schwing, and Kirillov 2021; Cheng et al. 2022; Ding et al. 2023). Recent work has explored diffusion models for segmentation (Amit et al. 2021; Baranchuk et al. 2021; Wang et al. 2024a; Gu, Chen, and Xu 2024), typically leveraging them as feature extractors within a discriminative framework. A key motivation behind SemFlow is the interpretation that diffusion models struggle to align their stochastic nature with the deterministic requirements of semantic segmentation. For this reason, SemFlow introduces rectified flows. In contrast, SymmFlow embraces the probabilistic nature of segmentation, accounting for inter-observer variability.
Semantic Image Generation
@@
Semantic image generation, the inverse of semantic segmentation, focuses on generating realistic images from semantic layouts (Li, Zhang, and Malik 2019; Zhu et al. 2020; Liu et al. 2019; Lv et al. 2022). Existing approaches generally fall into two categories. (1) GAN-based models (Zhu et al. 2017; Isola et al. 2017; Wang et al. 2018), although many of these methods struggle with mode collapse and produce only unimodal outputs. (2) Diffusion models, which treat semantic generation as a conditional generation task where semantic masks act as control signals (Wang et al. 2022). Some methods further integrate additional conditioning to enhance coherence (Zhang, Rao, and Agrawala 2023). However, these approaches often adopt asymmetric architectures with unidirectional generators, thereby complicating the unification of semantic segmentation and image generation.
Symmetrical Flow Matching
@@
Symmetrical Flow Matching (SymmFlow) unifies semantic segmentation and semantic synthesis as opposing flow processes, as illustrated in Figure 1. Given a data distribution $X$ (e.g. images) and a semantic representation $Y$ (e.g. masks or class labels), SymmFlow models bi-directional flows between them. The forward process transforms $X$ from noise, while simultaneously evolving $Y$ towards a noise-corrupted state. The reverse process inverts these transitions, allowing for the generation of $Y$ from $X$. Crucially, $Y$ is not restricted to having the same dimensionality as $X$, enabling flexible conditioning, such as global class labels for classification. This symmetrical formulation ensures sufficient entropy for image generation while preserving the semantic structure, making SymmFlow a generalizable framework for both segmentation and synthesis. The training procedure and objective are formalized in the section titled Training Objective. Techniques for obtaining segmentation and classification predictions using the proposed SymmFlow model are discussed in Classification and Segmentation. The importance of label dequantization for stable training is examined in Dequantization, and the complete framework is validated using a synthetic example presented in Toy Example.
Training Objective
@@
[…]
This eliminates the need for repeated evaluations across all possible class embeddings, significantly reducing inference time and computational cost. Additionally, the same process can be used to predict segmentation masks. The predicted class is determined as the closest label to the average of the model’s predictions. For segmentation, the class of each pixel is assigned based on the closest predefined class RGB code to the predicted pixel RGB value. This mapping is further explained in the Toy Example section.
Dequantization
@@
Toy Example
To illustrate the principles of Symmetrical Flow Matching, we consider a toy example where two classes form the interleaved, nonlinear structures in Figure 3.
Figure 3: Visual representation of the two-spiral dataset.
We train a multilayer perceptron (MLP) to model the joint evolution of data points $X$ and their class representation $Y$ under Symmetrical Flow Matching. The input to the model consists of the point coordinates $X$, a quantized encoding of their class $Y$ where Class A follows a uniform distribution in [-1.5, -0.5] and Class B in [+0.5, +1.5] — and a time variable $t$ ranging from 0 to 1. The learned flow is then used to sample trajectories via an Euler ODE solver with 20 timesteps from $t=0$ to $t=1$, reconstructing the underlying structure of the data. The resulting distributions, shown in Figure 4, demonstrate the ability of our framework to model structured transformations while maintaining class separability.
Figure 4: Visual representations of generated samples from the forward process of the model.
@@
Datasets
Semantic Segmentation and Generation: The model is evaluated using COCO-Stuff (Caesar, Uijlings, and Ferrari 2018) and CelebAMask-HQ (Lee et al. 2020), which contain 171 and 19 classes, respectively. Images and semantic masks are resized and cropped into 512$\times$512 pixels. A detailed analysis of the datasets is provided in the Appendix. Classification: We evaluate our model on MNIST (Deng 2012) and CIFAR-10 (Krizhevsky, Nair, and Hinton 2010) for classification, leveraging these low-resolution datasets to assess fundamental capabilities.
Metrics
Semantic Segmentation and Generation: For semantic segmentation, we evaluate with mean intersection over union (mIoU). For semantic image synthesis, we assess with the Fréchet inception distance (FID) (Heusel et al. 2017) and learned perceptual image patch similarity (LPIPS) (Zhang et al. 2018). Classification: The model is evaluated using classification accuracy.
Qualitative Evaluation
To complement the quantitative analysis, qualitative results are presented for classification, semantic segmentation, and image generation. For classification, generated samples from MNIST and CIFAR-10 are visualized in the Appendix to evaluate the diversity and fidelity of the synthesized outputs. For semantic segmentation, predicted masks on CelebAMask-HQ and COCO-stuff are compared against ground-truth annotations. For image generation, representative samples synthesized from CelebAMask-HQ and COCO-stuff are provided to illustrate visual quality.
Impact of Inference Steps
Baseline Comparison
Semantic Segmentation and Generation: For semantic segmentation on CelebAMask-HQ, DML-CSR (Zheng et al. 2022) and SegFace (Narayan, Vs, and Patel 2025) are adopted as baselines. On COCO-Stuff, DeeplabV2 (Caesar, Uijlings, and Ferrari 2018), MaskFormer (Cheng, Schwing, and Kirillov 2021), and SegFormer (Xie et al. 2021) serve as representative methods. For semantic image synthesis, the evaluation includes pix2pixHD (Wang et al. 2018), SPADE (Park et al. 2019), SC-GAN (Wang et al. 2021), BBDM (Li et al. 2023b), SDM (Wang et al. 2022), SCDM (Ko et al. 2024), and SCP-Diff (Gao et al. 2024). As a unified model capable of both segmentation and synthesis, SemFlow (Wang et al. 2024b) is also evaluated. Since official checkpoints and semantic segmentation evaluation scripts for SemFlow were not publicly available, the model was retrained and results were recomputed. The scripts and checkpoints are included in SymmFlow’s repository. Classification: For the classification task on CIFAR-10, the proposed model is compared to the Diffusion Classifier (Li et al. 2023a). Additionally, the quality of the generated images is assessed by benchmarking against a standard FM setup.
Implementation Details
For pixel-level implementations (e.g., classification), the U-Net architecture introduced in Guided Diffusion (Dhariwal and Nichol 2021) is employed. For latent-space models, the pre-trained VAE from Stable Diffusion is used for image encoding and decoding, alongside the U-Net backbone from Stable Diffusion 2.1 (Rombach et al. 2022). To ensure compatibility with SymmFlow, the number of input channels in the first layer and output channels in the final layer of the U-Net are doubled. Further details on hyperparameters and computational resources are provided in the Appendix.
Results & Discussion
Table 2: Performance comparison of benchmark solutions on semantic segmentation (SS) and semantic image synthesis (SIS) tasks across both COCO-Stuff and CelebAMask-HQ datasets. The number of steps indicates the number of functions applied to obtain the results. Legend:  Results were recomputed due to absence in the paper.
@@
SS DML-CSR (Zheng et al. 2022) 1 77.8
SegFace (Narayan, Vs, and Patel 2025) 1 81.6
DeeplabV2 (Caesar, Uijlings, and Ferrari 2018) 1 33.2
MaskFormer (Cheng, Schwing, and Kirillov 2021) 1 37.1
SegFormer (Xie et al. 2021) 1 46.7
SIS pix2pixHD (Wang et al. 2018) 1 54.7 / 0.529 111.5 / —
SPADE (Park et al. 2019) 1 42.2 / 0.487 33.9 / —
SC-GAN (Wang et al. 2021) 1 19.2 / 0.395 18.1 / —
BBDM (Li et al. 2023b) 200 21.4 / 0.370
SDM (Wang et al. 2022) 1000 18.8 / 0.422 15.9 / 0.518
SCDM (Ko et al. 2024) 250 17.4 / 0.418 15.3 / 0.519
SCP-Diff (Gao et al. 2024) 800 11.3 / —
Both SemFlow (Wang et al. 2024b) 25 *69.4 *35.7 32.6 / 0.393 *90.0 / 0.685
@@
Semantic Segmentation and Generation
Table 2 reports semantic segmentation and image synthesis results across benchmarks. SymmFlow consistently outperforms prior models on the synthesis task, achieving the lowest FID scores across datasets. Visualizations in Figure 5 confirm that the model produces high-fidelity, mask-consistent samples on both CelebAMask-HQ and COCO-Stuff, capturing structural details with strong adherence to the conditioning masks. While LPIPS is commonly used to assess perceptual diversity, it can be misleading in isolation—higher LPIPS values may result from poor image quality rather than true variability. Conversely, low LPIPS may indicate mode collapse or data leakage rather than accurate diversity. Therefore, interpreting LPIPS jointly with FID is essential. SymmFlow demonstrates the most favorable trade-off between image quality and diversity, reflecting strong generative capability without sacrificing semantic alignment.
In the segmentation task, SymmFlow achieves competitive performance compared to specialized segmentation baselines, particularly on COCO-Stuff. As shown in Figure 6, the model demonstrates semantic understanding beyond the ground-truth annotations; for instance, correctly identifying a laptop absent from the label map. However, performance is limited by the low-resolution latent representation (64$\times$64$\times$4), which hampers fine-grained accuracy. This is especially evident for small-area classes such as earrings or partially occluded features like eyebrows and ears, where segmentation quality deteriorates due to insufficient spatial detail. Despite this, the global segmentation structure remains coherent. SemFlow, while also achieving strong performance on the segmentation task, falls short on semantic image synthesis. Quantitative results indicate limited visual fidelity underscoring the limitations of the architecture in jointly modeling segmentation and generation without further design or training adjustments.
SymmFlow achieves this performance using only 25 function evaluations, whereas most diffusion-based methods require hundreds of denoising steps, resulting in significantly improved inference efficiency.

### v2

@@ Abstract
Flow Matching has emerged as a powerful framework for learning continuous transformations between distributions, enabling high-fidelity generative modeling. This work introduces Symmetrical Flow Matching (SymmFlow), a new formulation that unifies semantic segmentation, classification, and image generation within a single model. Using a symmetric learning objective, SymmFlow models forward and reverse transformations jointly, ensuring bi-directional consistency, while preserving sufficient entropy for generative diversity. A new training objective is introduced to explicitly retain semantic information across flows, featuring efficient sampling while preserving semantic structure, allowing for one-step segmentation and classification without iterative refinement. Unlike previous approaches that impose strict one-to-one mapping between masks and images, SymmFlow generalizes to flexible conditioning, supporting both pixel-level and image-level class labels. Experimental results on various benchmarks demonstrate that SymmFlow achieves state-of-the-art performance on semantic image synthesis, obtaining FID scores of 11.9 on CelebAMask-HQ and 7.0 on COCO-Stuff with only 25 inference steps. Additionally, it delivers competitive results on semantic segmentation and shows promising capabilities in classification tasks.
Codehttps://github.com/caetas/SymmetricFlow
1 Introduction
@@
Comprehending semantic content is a key challenge in computer vision. Classification (krizhevsky2012imagenet; he2016deep; dosovitskiy2020image) and segmentation (ronneberger2015u; xie2021segformer; cheng2022masked) allow models to analyze and structure images, while generative modeling enables the synthesis of new content (radford2015unsupervised; ho2020denoising; rombach2022high). Ideally, a unified framework would bridge these tasks, allowing models to both interpret and generate images in a two-way manner. It is conceivable that the ability to accurately comprehend and disentangle visual structures facilitates the generation of more semantically coherent and visually realistic images. Conversely, strong generative capabilities may aid in learning more expressive representations of images, as generating plausible content requires an implicit understanding of object relationships, textures, and context (he2022masked). These forward and backward relationships suggest that advances in one direction could naturally benefit the other, which fuels the pursuit of models that integrate both understanding and synthesis within a single cohesive framework.
@@
Most existing vision-only approaches treat these tasks individually. For classification, the extracted features are used in fully connected layers to obtain probability predictions for each class, while for segmentation, the features are used to train decoders for dense predictions. On the other hand, generative models, such as generative adversarial networks (GANs) (karras2021alias), diffusion models (karras2022elucidating), score matching models (song2020score), and Flow Matching models (lipman2022flow), synthesize images from a prior distribution. Although recent work has explored diffusion models for classification (li2023your) and segmentation (wu2024medsegdiffv2), these adaptations introduce significant limitations: classification is slow due to the need for iterative sampling across all possible classes, and segmentation frameworks are restricted to generating masks, lacking the ability to map back to realistic images.
@@
Recent works, such as SemFlow (wang2024semflow) and DepthFM (gui2024depthfm), have sought to unify image generation with semantic segmentation and depth estimation within a single generative framework. However, these models still suffer from the following key limitations: (1) they do not perform classification, restricting their applicability; (2) the image quality remains inferior to that of purely generative models; and (3) the models enforce a strict one-to-one mapping between segmentation or depth masks and images, requiring them to have the same number of channels, which limits their flexibility.
@@
2 Related Work
2.1 Flow Matching
@@
Flow Matching (FM) (lipman2022flow; liu2022flow) is a generative modeling framework that learns a velocity field to transform a source distribution into a target distribution through a continuous flow. By parameterizing this field with a neural network, the model becomes a Neural ODE (chen2018neural), allowing efficient sampling via numerical integration. It has advanced state-of-the-art performance in diverse applications, including image (esser2024scaling) and audio generation (vyas2023audiobox), as well as protein modeling (huguet2024sequence) and robotics (black2024pi_0). FM generalizes Continuous Normalizing Flows (CNFs) (chen2018neural; grathwohl2018ffjord) by eliminating the need for simulation during training, thereby making it computationally efficient. It also provides a unifying perspective on generative modeling, encompassing diffusion models, which can be seen as a special case where the probability path is defined via stochastic differential equations (lipman2024flow).
2.2 Generative Classifiers
@@
Several seminal works (hinton2007recognize; ranzato2011deep) have emphasized the importance of modeling the data distribution to enhance discriminative feature learning. Early approaches trained deep belief networks (hinton2006fast) to encode image data as latent representations, which were then used for recognition tasks. More recent advances in generative modeling have demonstrated the ability to learn efficient representations for global prediction tasks (he2022masked). In addition, generative models have been shown to improve adversarial robustness and calibration (huang2020neural). However, most prior work either jointly trains generative and discriminative models, or fine-tunes generative representations for downstream tasks. Diffusion Classifier (li2023your) specifically investigates the effectiveness of using diffusion models as image classifiers, albeit with severe inference constraints.
2.3 Semantic Segmentation
@@
Semantic segmentation aims to assign a semantic label to each pixel in an image. Conventional approaches rely on discriminative models, combining a strong feature extraction backbone with a task-specific decoder head for mask prediction (cheng2021per; cheng2022masked; ding2023mevis). Recent work has explored diffusion models for segmentation (amit2021segdiff; wang2024explore; gu2024diffusioninst), typically leveraging them as feature extractors within a discriminative framework. A key motivation behind SemFlow is the interpretation that diffusion models struggle to align their stochastic nature with the deterministic requirements of semantic segmentation. For this reason, SemFlow introduces rectified flows. In contrast, SymmFlow embraces the probabilistic nature of segmentation, accounting for inter-observer variability.
2.4 Semantic Image Generation
@@
Semantic image generation, the inverse of semantic segmentation, focuses on generating realistic images from semantic layouts (li2019diverse; liu2019learning; lv2022semantic). Existing approaches generally fall into two categories. (1) GAN-based models (zhu2017unpaired; isola2017image; wang2018high), although many of these methods struggle with mode collapse and produce only unimodal outputs. (2) Diffusion models, which treat semantic generation as a conditional generation task where semantic masks act as control signals (wang2022semantic). Some methods further integrate additional conditioning to enhance coherence (zhang2023adding). However, these approaches often adopt asymmetric architectures with unidirectional generators, thereby complicating the unification of semantic segmentation and image generation.
3 Symmetrical Flow Matching
@@
Symmetrical Flow Matching (SymmFlow) unifies semantic segmentation and semantic synthesis as opposing flow processes, as illustrated in Figure 1. Given a data distribution $X$ (e.g. images) and a semantic representation $Y$ (e.g. masks or class labels), SymmFlow models bi-directional flows between them. The forward process transforms $X$ from noise, while simultaneously evolving $Y$ towards a noise-corrupted state. The reverse process inverts these transitions, allowing for the generation of $Y$ from $X$. Crucially, $Y$ is not restricted to having the same dimensionality as $X$, enabling flexible conditioning, such as global class labels for classification. This symmetrical formulation ensures sufficient entropy for image generation while preserving the semantic structure, making SymmFlow a generalizable framework for both segmentation and synthesis. The training procedure and objective are formalized in Section 3.1. Techniques for obtaining segmentation and classification predictions using the proposed SymmFlow model are discussed in Section 3.2. The importance of label dequantization for stable training is examined in Section 3.3, and the complete framework is validated using a synthetic example presented in Section 3.4.
3.1 Training Objective
@@ 3.2 Classification and Segmentation
$y_{0}=y_{1}+\int_{1}^{0}v_{\theta}(x_{t},y_{t},t)_{y}dt.$ (8)
This eliminates the need for repeated evaluations across all possible class embeddings, significantly reducing inference time and computational cost. Additionally, the same process can be used to predict segmentation masks. The predicted class is determined as the closest label to the average of the model’s predictions. For segmentation, the class of each pixel is assigned based on the closest predefined class RGB code to the predicted pixel RGB value. This mapping is further explained in Section 3.4.
3.3 Dequantization
@@
3.4 Toy Example
To illustrate the principles of Symmetrical Flow Matching, we consider a toy example where two classes form the interleaved, nonlinear structures in Figure 3(a).
(a) Original spiral dataset.
(b) Generated samples.
Figure 3: Visualization of the spiral dataset and samples generated by our model.
We train a multilayer perceptron (MLP) to model the joint evolution of data points $X$ and their class representation $Y$ under Symmetrical Flow Matching. The input to the model consists of the point coordinates $X$, a quantized encoding of their class $Y$ - where Class A follows a uniform distribution in [-1.5, -0.5] and Class B in [+0.5, +1.5] — and a time variable $t$ ranging from 0 to 1. The learned flow is then used to sample trajectories via an Euler ODE solver with 20 timesteps from $t=0$ to $t=1$, reconstructing the underlying structure of the data. The resulting distributions, shown in Figure 3(b), demonstrate the ability of our framework to model structured transformations while maintaining class separability.
@@ 4 Experiments
Table 2: Performance comparison of benchmark solutions on semantic segmentation ($S$) and semantic image synthesis (SIS) tasks across both COCO-Stuff and CelebAMask-HQ datasets. The number of steps indicates the number of functions applied to obtain the results. Legend:  Recomputed by us; — Not available; ✗ Not applicable (method cannot perform the task).
@@
$S$ DML-CSR (zheng2022decoupled) 1 77.8
SegFace (narayan2025segface) 1 81.6
DeeplabV2 (caesar2018coco) 1 33.2
MaskFormer (cheng2021per) 1 37.1
SegFormer (xie2021segformer) 1 46.7
SIS pix2pixHD (wang2018high) 1 54.7 / 0.529 111.5 / —
SPADE (park2019semantic) 1 42.2 / 0.487 33.9 / —
SC-GAN (wang2021image) 1 19.2 / 0.395 18.1 / —
BBDM (li2023bbdm) 200 21.4 / 0.370
ControlNet (zhang2023adding) 20 24.0 / 0.528 36.6 / 0.671
SDM (wang2022semantic) 1000 18.8 / 0.422 15.9 / 0.518
SCDM (ko2024stochastic) 250 17.4 / 0.418 15.3 / 0.519
SCP-Diff (gao2024scp) 800 11.3 / —
Both SemFlow (wang2024semflow) 25 *69.4 *35.7 32.6 / 0.393 *90.0 / 0.685
@@
4.1 Datasets
Semantic Segmentation and Generation: The model is evaluated using COCO-Stuff (caesar2018coco) and CelebAMask-HQ (lee2020maskgan), which contain 171 and 19 classes, respectively. Images and semantic masks are resized and cropped into $\times512$ pixels. A detailed analysis of the datasets is provided in Appendix A. Classification: We evaluate our model on MNIST (deng2012mnist) and CIFAR-10 (krizhevsky2010cifar) for classification, leveraging these low-resolution datasets to assess fundamental capabilities.
4.2 Metrics
Semantic Segmentation and Generation: For semantic segmentation, we evaluate with mean intersection over union (mIoU). For semantic image synthesis, we assess with the Fréchet inception distance (FID) and learned perceptual image patch similarity (LPIPS). Classification: The model is evaluated using classification accuracy. Additional information is available in Appendix D.
4.3 Qualitative Evaluation
To complement the quantitative analysis, qualitative results are presented for classification, semantic segmentation, and image generation. For classification, generated samples from MNIST and CIFAR-10 are visualized in Appendix G to evaluate the diversity and fidelity of the synthesized outputs. For semantic segmentation, predicted masks on CelebAMask-HQ and COCO-stuff are compared against ground-truth annotations. For image generation, representative samples synthesized from CelebAMask-HQ and COCO-stuff are provided to illustrate visual quality.
4.4 Impact of Inference Steps
4.5 Baseline Comparison
Semantic Segmentation and Generation: For semantic segmentation on CelebAMask-HQ, DML-CSR (zheng2022decoupled) and SegFace (narayan2025segface) are adopted as baselines. On COCO-Stuff, DeeplabV2 (caesar2018coco), MaskFormer (cheng2021per), and SegFormer (xie2021segformer) serve as representative methods. For semantic image synthesis, the evaluation includes pix2pixHD (wang2018high), SPADE (park2019semantic), SC-GAN (wang2021image), BBDM (li2023bbdm), ControlNet (zhang2023adding), SDM (wang2022semantic), SCDM (ko2024stochastic), and SCP-Diff (gao2024scp). SemFlow (wang2024semflow), a prior model aiming to unify both segmentation and synthesis, is also evaluated. Since official checkpoints and semantic segmentation evaluation scripts for SemFlow were not publicly available, the model was retrained and results were recomputed. The scripts are included in SymmFlow’s repository for reproducibility. Classification: For the classification task on CIFAR-10, the proposed model is compared to the Diffusion Classifier (li2023your). Additionally, the quality of the generated images is assessed by benchmarking against a standard FM setup.
4.6 Implementation Details
For pixel-level implementations (e.g., classification), the U-Net architecture introduced in Guided Diffusion (dhariwal2021diffusion) is employed. For latent-space models, the pre-trained VAE from Stable Diffusion (rombach2022high) is used for image encoding and decoding, alongside the U-Net backbone from Stable Diffusion 2.1. To ensure compatibility with SymmFlow, the number of input channels in the first layer and output channels in the final layer of the U-Net are doubled. Further details on hyperparameters and computational resources are provided in Appendix B.